Skip is now three times as old as Paul. In 5 years, Skip will be twice as old as Paul will be then. What will their ages be in five years?
Answer by mananth(16946)

let paul's age be x
skip's age = 3x
After 5 years their ages will be
x+5 and 3x+5
2(x+5)= 3x+5
2x+10 = 3x+5
3x-2x=10-5
x=5 which is paul's age
skip's age is 3 times paul's age = 15 years